1.Suppose you kick a soccer ball with an initial speed of 12 m/s at a certain upward angle. After 4.0 seconds, the ball covers a 35-m horizontal displacement. What is the angle of inclination of your initial velocity? Use tan-1 in solving for the answer
1
Expert's answer
2020-12-10T11:03:28-0500
"x=v_xt=vt\cos{\theta}\\
\theta=\cos^{-1}{\frac{x}{vt}}\\
\theta=\cos^{-1}{\frac{35}{4\cdot12}}=43.18\degree"